A ball initially at rest rolls down a hill with an acceleration of 4.5 m/s^2
If it accelerates for 9 seconds, how far will it move?
1
Expert's answer
2018-09-04T11:04:08-0400
S=v_0 t+(at^2)/2, v_0=0 → S=(at^2)/2=(4.5∙9^2)/2=182.25 m.